# Regularized second-order optimization of tensor-network Born machines

### v1

@@ IV Results
We use these two dataset to compare the different optimization algorithms discussed in this work- steepest descent, Newton’s method, and the regularized Newton’s method by smoothing of the landscape. For both the BAS and the MNIST datasets, we prepare datasets of images with 7 by 7 grid. For the MNIST dataset, we course-grain the original grid of 28 by 28 pixels to 7 by 7 pixels and set a threshold to convert it from grayscales to a binary image. We represent the 2-dimensional grid of 49 pixels using a 1-dimensional MPS, unraveling the pixels for the MNIST dataset in a snake-like configuration to retain as much of the original notion of locality as possible in the transition from 2 to 1 dimensions.
@@
Implementing our new optimization algorithm, we define a regularization schedule that varies the regularization constant throughout the optimization. By starting with relatively large values of $\epsilon$ that decrease with each iteration, we avoid getting stuck in local minima early at initial iterations, while maintaining good optimization accuracy later when the optimizer reaches the proximity of the desired global minimum. In our numerical experiments, we chose an exponentially decreasing regularization schedule, starting from a moderate value of $\epsilon=0.025$ and scaling down exponentially for later sweeps.
@@
Figure 4: A comparison of loss curves for training of TNBM on the bars and stripes (BAS) and the MNIST datasets using different optimization algorithms. We compare our method against the steepest descent and the non-regularized Newton’s method. The figure presents the loss as a function of the number of iterations, while the bands represent the standard deviation of the loss curves between five randomly initialized realizations. The bottom row presents a zoom-in view on the region of lower loss values reached after a single forward and backward sweep.
@@
For both datasets, we set the tensor network as a 49-sites MPS with a bond dimension of $\chi=5$, and optimize it using 5 forward and backward sweeps. We train our model using 100 samples from the dataset, and average the results over 5 random initialization of the model. For each of these optimization algorithms, we iterate across all tensors back and forth, taking a single optimization step per tensor. Figure 4 presents the resulting loss curves of the TNBM model as a function of the number of iterations, for both the BAS and the MNIST datasets. In the figure we compare the loss curve of our regularized optimization algorithm with simple gradient descent and with an unaltered version of Newton’s method. We observe that the regularization mechanism improves the quality of the converged model, reaching lower loss values faster compared to gradient descent or the naive newton optimizer.
@@
The loss curve reveals different stages of optimization of the regularized Newton optimization method. In the first sweep, for about 50 iteration, the loss curve of the regularized optimization follows the same trend of both Newton and gradient descent optimization. By the second sweeps the differences between optimizers start to emerge, with the unaltered constrained Newton’s method shows an improvement over both gradient descent and the regularized Newton. At this stage the regularization increases the NLL score of the model, as the optimizer ignores the effect of nearby singularities. As the regularization constant diminishes, the bump in the loss curve of the regularized optimization flattens, and the regularized optimizer reaches lower loss values at early iterations, compared to the other optimizers. There are several differences between the results of the two datasets. For the BAS dataset, the optimization curve is smooth, with diminishing returns for subsequent sweeps. For the MNIST dataset, the optimization curve after the first sweep is structured as sequence of downward slopes interleaved with plateau sections. This jagged structure is due to the difference in pixel variance between peripheral and central pixels on the 2-dimensional grid. The comparison between the optimizers for the MNIST dataset shows a consistent, though less significant, advantage compared to the results with the bars and stripes dataset. The reason behind this difference is probably related to the complexity of the MNIST dataset, containing 10 different labels digit labels, as well as a more complex 2-dimensional geometric structure that is hard to capture using a 1-dimensional MPS. Those properties increase the difficulty of optimization for all models equally, which lead to closer results between the three of them.
@@
Figure 5: Loss curves of continuous embedding MPS for various optimization methods: simple gradient-descent, complex-valued gradient descent, Newton’s method optimization and Regularized newton on manifold
@@ V Conclusion and discussion
Table 1: Complexity comparison between the gradient descent and Newton’s method optimizers.
In this work we introduced an improved optimization algorithm for tensor-network Born machines, capable of reaching high quality convergence faster than simple gradient descent while avoiding local minima. The algorithm utilizes restricted version of the Newton optimization method on manifold, together with a regularization of the loss function landscape to mitigate singularities. We described two versions of regularization, introducing either smoothing or shifting of the logarithmic singularities, to help the optimizer avoid local minima and converge to a lower loss point. We have shown that this combination robustly find better quality minimal solutions faster than gradient descent, demonstrating the effect for the discrete datasets of bars-and-stripes and MNIST, and for the continuous IRIS dataset with a continuous TNBM. Our algorithm can be naturally expanded and implemented for various tensor-network architectures, such as 2-dimensional PEPS and the tree-like MERA network, which can improve model performances for datasets with different geometrical properties.
The effect of regularization on the landscape mimics the effect of complex-valued tensors by effectively getting around the logarithmic pole with a constant imaginary shift of the overlap, while avoiding the memory and computational overhead of handling complex numbers. This approach restores the feasibility of using real-valued tensors for optimization, enabling efficient training without the need for explicitly introducing complex-valued representations.
A potential concern with our method is the computational cost of Newton’s optimization, particularly the need to calculate the Hessian. However, for the case of TNBM, we can use the structure of loss function to avoid explicitly computing and storing the full Hessian matrix. Instead, we use iterative solvers that only require Hessian-vector products, which are efficient to compute, significantly improving efficiency. For the NLL loss function, both single-tensor gradient and Hessian calculations involve only $O(DN_{\rm s})$ operations, for tensor dimension $D$ and $N_{\rm s}$ samples, with memory complexity of $O(D)$ for gradients and $O(D^{2})$ if storing the full dense Hessian were necessary. Similarly, Newton’s optimization step can be computed by either a direct matrix inversion or by using iterative methods.
@@
It should be noted that the regularization mechanism does not hold as a general-purpose solution for local minima trapping. The general task of training a generative model using limited resources is a challenging task, and naturally, the quantum inspired approach taken by the TNBM can at times have difficulties to converge to a good solution due to various reasons. These reasons may include incompatibility of the chosen geometry and dimensions of the tensor network, and the lack of sufficient training data or computation time, as well as multi-body patterns that are hard for the model to recognize. Many times these problems manifest as convergence to local minima, often by effects distributed along many tensor cores, which cannot be solved by regularization. Despite these limitations, our optimization method provides significant acceleration in many cases, enhancing the efficiency of TNBMs for real-world applications.
@@ Appendix A Derivation of the regularized gradient and Hessian
$\frac{\partial}{\partial T}\text{grad}\,\overline{L}_{\epsilon}(T)=2\sum_{x}{p_{x}\frac{(T,w_{x})\left((T,w_{x})I+T\otimes w_{x}\right)}{{(T,w_{x})^{2}+\epsilon}}}+2\sum_{x}{p_{x}\frac{(T,w_{x})^{2}-\epsilon}{\left((T,w_{x})^{2}+\epsilon\right)^{2}}{\rm proj}_{T}(w_{x})\otimes w_{x}}$ (24)
@@
${\rm Hess}\,\overline{L}_{\epsilon}(T)=2\sum_{x}{p_{x}\frac{(T,w_{x})^{2}\left(I-T^{\otimes 2}\right)}{{(T,w_{x})^{2}+\epsilon}}}+2\sum_{x}{p_{x}\frac{(T,w_{x})^{2}-\epsilon}{\left((T,w_{x})^{2}+\epsilon\right)^{2}}\left({\rm proj}_{T}(w_{x})\right)^{\otimes 2}}$ (25)
@@
$\hat{H}_{\epsilon}^{\rm abs}=2\sum_{x}{p_{x}\frac{(T,w_{x})^{2}\left(I-T^{\otimes 2}\right)}{{(T,w_{x})^{2}+\epsilon}}}+2\sum_{x}{p_{x}\frac{\absolutevalue{(T,w_{x})^{2}-\epsilon}}{\left((T,w_{x})^{2}+\epsilon\right)^{2}}\left({\rm proj}_{T}(w_{x})\right)^{\otimes 2}}$ (26)

### v2

@@ IV Results
We use these two dataset to compare the different optimization algorithms discussed in this work- steepest descent, Newton’s method, and the regularized Newton’s method by smoothing of the landscape. For both the BAS and the MNIST datasets, we prepare datasets of images with 7 by 7 grid. For the MNIST dataset, we perform downsampling on the original 28 by 28 grid via average pooling, then binarizing the result using a threshold. We represent the final 2-dimensional grid of 49 binary pixels using a 1-dimensional MPS, unraveling the pixels for the MNIST dataset in a snake-like configuration to retain as much of the original notion of locality as possible in the transition from 2 to 1 dimensions. We then scramble all subgroups of different labels together, drawing random samples from arbitrary labels as the training set.
@@
Implementing our new optimization algorithm, we define a regularization schedule that varies the regularization constant throughout the optimization. By starting with relatively large values of $\epsilon$ that decrease with each iteration, we smooth out peaks and explore the whole landscape without barriers and avoid getting stuck in local minima early at initial iterations, while maintaining good optimization accuracy later when the optimizer reaches the proximity of the desired global minimum. In our numerical experiments, we chose an exponentially decreasing regularization schedule, starting from a moderate value of $\epsilon=0.025$ and scaling down exponentially for later sweeps.
@@
Figure 4: A comparison of loss curves for training of TNBM on the bars and stripes (BAS) and the MNIST datasets using different optimization algorithms. We compare the regularized Newton method (green) against the steepest descent (blue) and the vanilla Newton’s method (red). The figure presents the loss as a function of the number of iterations, while the bands represent the standard deviation of the loss curves between five randomly initialized realizations. The bottom row presents a zoom-in view on the region of lower loss values reached after a single forward and backward sweep.
@@
For both datasets, we set the tensor network as a 49-sites MPS with a bond dimension of $\chi=5$, and optimize it using 5 forward and backward sweeps. We train our model using 100 samples from the dataset, and average the results over 5 random initialization of the model. For each of these optimization algorithms, we iterate across all tensors back and forth, taking a single optimization step per tensor. Figure 4 presents the resulting loss curves of the TNBM model as a function of the number of iterations, for both the BAS and the MNIST datasets. In the figure we compare the loss curve of our regularized optimization algorithm with simple gradient descent and with an unaltered version of Newton’s method. The loss score metric is presented in its original form for all methods, i.e. them non-regularized loss as written in Eqs. 1 and 4. We observe that the regularization mechanism improves the quality of the converged model, reaching lower loss values faster compared to gradient descent or the naive newton optimizer.
@@
The loss curve reveals different stages of optimization of the regularized Newton optimization method. In the first sweep, for about 50 iteration, the loss curve of the regularized optimization follows the same trend of both Newton and gradient descent optimization. By the second sweep, the differences between optimizers start to emerge, with the unaltered constrained Newton’s method shows an improvement over both gradient descent and the regularized Newton. At this stage the regularization increases the NLL score of the model, as the optimizer ignores the effect of nearby singularities. As the regularization constant diminishes, the bump in the loss curve of the regularized optimization flattens, and the regularized optimizer reaches lower loss values at early iterations, compared to the other optimizers. There are several differences between the results of the two datasets. For the BAS dataset, the optimization curve is smooth, with diminishing returns for subsequent sweeps. For the MNIST dataset, the optimization curve after the first sweep is structured as a descending staircase- a sequence of downward slopes interleaved with plateau sections. This jagged structure is due to the difference in pixel variance between peripheral and central pixels along the snake-like path n the two dimensional grid.
The comparison between the optimizers for the MNIST dataset shows a consistent, though less significant, advantage compared to the results with the BAS dataset. The reason behind this difference is likely related to the complexity of the MNIST dataset, containing 10 different labels digit labels, as well as a more complex two dimensional geometric structure that is hard to capture using a one dimensional MPS. Those properties increase the difficulty of optimization for all models equally, which lead to closer results between the three of them.
@@
Figure 5: Loss curves of continuous embedding MPS for various optimization methods: simple gradient-descent, Newton’s method optimization and the two variants of regularized newton optimization on manifold, the smoothing regularization in Eq. 16 and bias regularization in in Eq. 19.
@@ V Conclusion and discussion
Table 1: Complexity comparison between gradient descent and Newton’s method optimizers, for calculation of gradients, Hessians and full iteration step.
In this work we introduced an improved optimization algorithm for tensor-network Born machines, capable of reaching high quality convergence faster than simple gradient descent while exploring the global manifold and avoiding local minima. The algorithm utilizes restricted version of the Newton optimization method on manifold, together with a regularization of the loss function landscape to mitigate singularities. We described two versions of regularization, introducing either smoothing or shifting of the logarithmic singularities, to help the optimizer avoid local minima and converge to a lower loss point. We have shown that this combination robustly find better quality minimal solutions faster than gradient descent, demonstrating the effect for the discrete datasets of bars-and-stripes and MNIST, and for the continuous IRIS dataset with a continuous TNBM. Our algorithm can be naturally expanded and implemented for various tensor-network architectures, such as two-dimensional PEPS [42, 15] and the tree-like MERA network [43, 44], which can improve model performances for datasets with different geometrical properties.
The effect of regularization on the landscape mimics the effect of complex-valued tensors by effectively getting around the logarithmic pole with a constant imaginary shift of the overlap, while avoiding the memory and computational overhead of handling complex matrices. This approach restores the feasibility of using real-valued tensors for optimization, enabling efficient training without the need for explicitly introducing complex-valued representations.
A potential concern with our method is the computational cost of Newton’s optimization, particularly the need to calculate the Hessian. However, for the case of TNBM, we can use the structure of loss function to avoid explicitly computing and storing the full Hessian matrix. Instead, we use iterative solvers that only require Hessian-vector products. Instead of constructing the Hessian matrix out explicitly causing huge computational and storage waste, it is more efficient to use iterative solvers for the Newton step, which only require evaluating the projection of the Hessians on vectors, significantly improving efficiency.For the NLL loss function, both single-tensor gradient and Hessian calculations involve only $O(DN_{\rm s})$ operations, for tensor dimension $D$ and $N_{\rm s}$ samples, with memory complexity of $O(D)$ for gradients and $O(D^{2})$ if storing the full dense Hessian were necessary. Similarly, Newton’s optimization step can be computed by either a direct matrix inversion or by using iterative methods.
@@
It should be noted that the regularization mechanism does not hold as a general-purpose solution for local minima trapping in TNBMs. The general task of training a generative model using limited resources is a challenging task, and naturally, the quantum inspired approach taken by the TNBM can at times have difficulties to converge to a good solution due to various reasons, such as incompatibility of the chosen geometry and dimensions of the tensor network, and the lack of sufficient training data or computation time, as well as multi-body patterns that are hard for the model to recognize. Many times these problems manifest as convergence to local minima, often by effects distributed along many tensor cores, which cannot be solved by regularization. Despite these limitations, our optimization method provides significant acceleration in many cases, enhancing the efficiency of TNBMs for real-world applications.
@@ Appendix A Derivation of the regularized gradient and Hessian
$\frac{\partial}{\partial T}\text{grad}\,\overline{L}_{\epsilon}(T)=2\sum_{x}{p_{x}\frac{(w_{x},T)\left((T,w_{x})I+T\otimes w_{x}\right)}{{|(T,w_{x})|^{2}+\epsilon}}}+2\sum_{x}{p_{x}\frac{|(T,w_{x})|^{2}-\epsilon}{\left(|(T,w_{x})|^{2}+\epsilon\right)^{2}}{\rm proj}_{T}(w_{x})\otimes w_{x}}$ (24)
@@
${\rm Hess}\,\overline{L}_{\epsilon}(T)=2\sum_{x}{p_{x}\frac{|(T,w_{x})|^{2}\left(I-T^{\otimes 2}\right)}{{|(T,w_{x})|^{2}+\epsilon}}}+2\sum_{x}{p_{x}\frac{|(T,w_{x})|^{2}-\epsilon}{\left(|(T,w_{x})|^{2}+\epsilon\right)^{2}}\left({\rm proj}_{T}(w_{x})\right)^{\otimes 2}}$ (25)
@@
[…]